OpenStudy (anonymous):

4 h minus 6 over 8 = -3

OpenStudy (anonymous):

\[\frac{ 4h-6 }{ 8 }=-3\] What would your very first step be? Remember to use your elementary school rules. BODMAS. Brackets first, then division, then multiplication, then Addition and last but not least, subtraction.

OpenStudy (kropot72):

Is it \[(1)\ \frac{4h-6}{8}=-3\] or \[(2)\ 4h-\frac{6}{8}=-3\]